how do determine the equation of a line using the information "the slope is -3 and it passes through (2,-1)
thank you

Answer by palanisamy(496)

Given, the slope of a line is -3 and it passes through (2,-1)
So the equation of the line is
(y+1) = -3(x-2)
y+1 = -3x+6
3x+y-5=0